a rectangular box is to be constructed with its length four times the size of its width, and its height one half the size of its length. the volume of the box must be 1000 cubic centimeters. find the dimensions of the box
Answer by ankor@dixie-net.com(22740)

a rectangular box is to be constructed with its length four times the size of its width, and its height one half the size of its length.
the volume of the box must be 1000 cubic centimeters. find the dimensions of the box
:
Let x = the width of the box
then
4x = the length of the box
and
2x = the height of the box (half the length
:
Volume: L*W*H
therefore
x*4x*2x = 1000
8x^3 = 1000
Find the cube root of both sides
2x = 10
x = 5 cm is the width
then
4(5) = 20 cm is the length
and
2(5) = 10 cm = the height
:
:
Check this: 5 * 20 * 10 = 1000